OpenStudy (anonymous):

3n^3-4n^2+9n-12 factor by grouping

OpenStudy (anonymous):

3n^3+9n = 3n(n^2+3) -4n^2-12 = -4(n^2-3) So, you get (n^2-3)(3n-4)

OpenStudy (anonymous):

n^2(3n-4) 3(3n-4) so our factor are n^2 3 and 3n

OpenStudy (anonymous):

\[rewrite \to 3n ^{3}+9n-4n ^{2}-12\] \[3n \left( n ^{2}+3 \right)-4\left( n ^{2}+3 \right)\]

OpenStudy (anonymous):

I mistyped n^2-3 instead of n^2+3. Can't copy paste and fix with this tool easily.

OpenStudy (anonymous):

It should be as follows: 3n^3+9n = 3n(n^2+3) -4n^2-12 = -4(n^2+3) So, you get (n^2+3)(3n-4)